# Comparing Machine Learning Algorithms by Union-Free Generic Depth

### 20 Dec 2023 v1

@@ 1.1 Motivation
Of course, a descriptive analysis of samples of partial orders requires a completely different mathematical apparatus than the analysis of standard data. A suitable formal framework is by no means obvious. [^8] Many approaches proposed for analyzing poset-valued data rely on distance measures, see, e.g., Critchlow (1985); Brandenburg et al. (2012). For further discussion see Blocher et al. (2022), Section 2. Fortunately, it turns out that the concept of (data) depth function and the data representation given by formal concept analysis can be promisingly combined and applied to poset-valued random variables. [^9] The representation of posets via formal concept analysis can be expressed as a closure operator, see Definition 1. Therefore, formal concept analysis is not necessary to follow this paper. For more information and discussion of this representation, see Blocher et al. (2022). In this paper, we adapt the concept of depth to poset-valued random variables by introducing the union-free generic (ufg) depth. In general, depth functions define a notion of centrality and outlyingness of observations with respect to an entire data cloud or an underlying distribution, see e.g. Serfling and Zuo (2000); Mosler (2002). So far, depth functions have mainly been applied to $\mathbb{R}^{d}$-valued random variables. Exceptions are e.g. the definition of depth functions on lattices in Schollmeyer (2017b, a), on total orders, see Goibert et al. (2022), and a general discussion on depth functions based on formal concept analysis in Blocher and Schollmeyer (2023). Since we transfer this depth concept to poset-valued data, some classical descriptive statistics can naturally be adapted to this particular non-standard data type.
@@ 1.2 Comparing Machine Learning Algorithms
(I) First, we compare algorithms not with respect to one unidimensional criterion like, e.g., balanced accuracy, but instead look at a whole set of performance measures. We then judge one algorithm as at least as good as another one if it is not outperformed with respect to any of these performance measures. With this, for every data set, we get a partial order of algorithms and since we are not looking at only one, but a whole population/sample of data sets, we get a poset-valued random variable. Importantly, we do not hold the view that there is an underlying true (random) total order together with a coarsening mechanism that generates the (random) partial order. Such views are often termed epistemic, cf., Couso and Dubois (2014). Applications of this view in the context of partial order data can be found for example in Lebanon and Mao (2008); Nakamura et al. (2019). In contrast, within the nomenclature of Couso and Dubois (2014), we see our approach more in the spirit of the ontic view that is usually applied to set-valued data. [^10] Concrete applications can be found e.g. in Plass et al. (2015b) for the ontic, and in Plass et al. (2015a) for the epistemic view. A case where the ontic and the epistemic view coincide is discussed in Schollmeyer (2019). Beyond this, in the field of partial identification in the context of generalizing confidence intervals to confidence regions for the so-called identified set, the question about an ontic vs. an epistemic view is also implicitly asked (without reference to these terms) by asking if such a confidence set should cover the true parameter or the whole identified set with prespecified probability, cf., Stoye (2009). However, since in our case the random objects are partial orders, the term ontic seems to fit not perfectly. Instead, we understand poset-valued data as a special type of non-standard data. [^11] Note that we do not want to generally rule out epistemic treatment, but this is not the focus of this paper. Such a treatment could use that every partial order can be described by the set of all its linear extensions. For further discussions, c.f., Blocher et al. (2022).
@@ 1.3 Outline of the Paper
This paper is an extended version of Blocher et al. (2023) presented at the 13th International Symposium on Imprecise Probabilities: Theories and Applications (ISIPTA 2023) conference in Oviedo, Spain. Compared to the ISIPTA 2023 conference, this paper focuses more on the application and comparison with other methods for evaluating poset-valued data, see Section 6. We also extend the analysis and implementation of the ufg depth, see Sections 4 and 5.
@@
Our paper is organized as follows: In Section 2, we briefly discuss the required mathematical definitions and concepts. We give a formal definition of our depth function, the ufg depth, in Section 3 and discuss some of its properties in Section 4. While Section 5 prepares our application by providing implementation details, Section 6 is devoted to applying our framework to specific examples, namely the analysis of the goodness of classification algorithms on different data sets. First, we focus on comparing our benchmarking approach to further benchmarking concepts, see Section 6.1 and afterward, we give a detailed demonstration of the ufg method, see Section 6.2. Section 7 concludes by elaborating on some promising perspectives for future research.
@@ 2 Preliminaries
The concept for analyzing poset-valued observations presented in this paper is based on a closure operator on $\mathcal{P}$, see Section 3. In general, a closure operator $\gamma_{\Omega}:2^{\Omega}\to 2^{\Omega}$ on set ${\Omega}$ is an operator which is extensive (for $A\subseteq{\Omega},\>A\subseteq\gamma_{\Omega}(A)$ holds), increasing, (for $A,B\subseteq{\Omega}$ with $A\subseteq B,\>\gamma_{\Omega}(A)\subseteq\gamma_{\Omega}(B)$ is true) and idempotent (for $A\subseteq{\Omega},\;\gamma_{\Omega}(A)=\gamma_{\Omega}(\gamma_{\Omega}(A))$ holds). The set $\gamma_{\Omega}(2^{\Omega})$ is called the closure system and we say that $\gamma(B)$ for $B\subseteq\Omega$ is the closure of $B$. In what follows, we use that every closure operator (and therefore the closure system) can be uniquely described by an implicational system. Generally, an implicational system $\mathcal{I}$ on ${\Omega}$ is defined by a subset of $2^{\Omega}\times 2^{\Omega}$. The implicational system corresponding to the closure operator $\gamma_{\Omega}$ is defined by all pairs $(A,B)\in 2^{\Omega}\times 2^{\Omega}$ satisfying $\gamma_{\Omega}(A)\supseteq\gamma_{\Omega}(B)$. For short, we denote this by $A\to B$. For more details on closure operators and implicational systems, see Bertet et al. (2018).
We aim to define a centrality and outlyingness measure on the set of all posets $\mathcal{P}$ based on a fixed and finite set $M$. In general, functions that measure the centrality of a point with respect to an entire data cloud or an underlying distribution are called (data) depth functions. Depth functions on $\mathbb{R}^{d}$ have been studied intensively by, e.g., Serfling and Zuo (2000) and Mosler and Mozharovskyi (2022), and various notions of depth have been defined, such as Tukey’s depth, see Tukey (1975), and simplicial depth, see Liu (1990). The idea behind the ufg depth introduced here is an adaptation of the simplicial depth on $\mathbb{R}^{d}$ to posets. The simplicial depth on $\mathbb{R}^{d}$ is based on the convex hull/closure operator that is defined as follows:
@@
where $X_{1},\ldots,X_{d+1}\overset{iid}{\sim}\nu$ with iid being independent and identically distributed for short. When we consider a sample $x_{1},\ldots,x_{n}\in\mathbb{R}^{d};\>n\in\mathbb{N}$, we use the empirical probability measure instead of a probability measure $\nu$. Thus, for a sample $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$ with empirical measure $\nu_{n}$ we obtain as empirical simplicial depth
$D_{n}\colon\begin{array}[]{l}\mathbb{R}^{d}\to[0,1],\\ x\mapsto\binom{n}{d+1}\sum_{1\leq i_{1}<\ldots<i_{d+1}\leq n}1_{\gamma_{\mathbb{R}^{d}}\{x_{i_{1}},\ldots,x_{i_{d+1}}\}}(x).\end{array}$
@@ 3 Union-Free Generic Depth on Posets
Now, we introduce the union-free generic (ufg) depth function for posets which is in the spirit of the simplicial depth function, see Section 2. To define the depth function, we start, similar to the simplicial depth, by defining a closure operator on $\mathcal{P}$.
@@
This closure operator maps a set of posets $P$ onto the set of posets where each poset is a superset of the intersection of $P$ and a subset of the union of $P$. In other words, any $p$ lying in the closure of $\gamma(P)$ satisfies the following condition: First, every pair $(y_{1},y_{2})\in M\times M$ that lies in every poset in $P$ is also contained in $p$, and second, for every pair $(y_{1},y_{2})$ that lies in $p$, there exists at least one $\tilde{p}\in P$ such that $(y_{1},y_{2})\in\tilde{p}$. Note that while the intersection of posets defines a poset again, this does not hold for the union. Analogously to the definition of the simplicial depth, we now only consider a subset of $2^{\mathcal{P}}$.
@@
$\mathscr{S}$ is a proper subset of $2^{\mathcal{P}}$, see Theorem 3 for details, which reduces $2^{\mathcal{P}}$ by redundant elements in the following sense: First, all subsets $P\subseteq\mathcal{P}$ with $\gamma(P)=P$ are trivial and therefore not included in $\mathscr{S}$. Second, if there exists a proper subset $\tilde{P}\subsetneq P$ with $\gamma(\tilde{P})=\gamma(P)$, then $P$ is also not in $\mathscr{S}$. This follows by setting $\ell=1$ and $A_{1}=\tilde{P}$, which defines a family contradicting Condition (C2). These two properties can be generalized to arbitrary closure systems, and referring to Bastide et al. (2000), we call a set fulfilling these two properties generic. The final reduction is to delete also all sets $P$ where $P$ can be decomposed by a family of proper subsets $(A_{i})_{i\in\{1,\ldots,\ell\}}$ of $P$. Moreover, the union of $(\gamma(A_{i}))_{i\in\{1,\ldots,\ell\}}$ needs to equal $\gamma(P)$. Note that due to extensitivity, the assumption $\cup_{i\in\{1,\ldots,\ell\}}\gamma(A_{i})\subseteq\gamma(P)$ is always true. We call sets respecting this third part union-free. Thus $\mathscr{S}$ consists of elements that are union-free and generic.
@@
The empirical version of the ufg depth uses the empirical probability measure $\nu_{n}$ given by a sample of posets $\underline{p}=(p_{1},\ldots,p_{n}),\>n\in\mathbb{N}$ instead of the probability measure $\nu$ in Definition 3.
@@
The empirical ufg depth of a poset $p$ is therefore the normalized weighted sum of drawn sets $S\in\mathscr{S}$ which imply $p$. Note that when restricting $\mathscr{S}$ to the set $\mathscr{S}_{obs}:=\{S\in\mathscr{S}\mid\mbox{ all }p\in S\mbox{ are observed}\}$, this does not change the depth value. This holds since for other elements $S\in\mathscr{S}$, the empirical measure for at least one $p\in S$ is zero.
@@
Returning to Example 1, suppose that we observe $(p_{1},p_{2},p_{3})$. Then for the trivial poset $p_{\Delta}$, the empirical depth is $D_{n}(p_{\Delta})=1/2$. For the set $p_{4}$ given by the transitive hull of $\{(y_{3},y_{1})\}$, the value of the empirical depth is zero. For $p_{total}$ given by the transitive hull of $\{(y_{1},y_{3}),(y_{3},y_{2})\}$, the empirical depth value is again zero.

### 21 Feb 2024 v3

@@ 1.1 Motivation
Of course, a descriptive analysis of samples of partial orders requires a completely different mathematical apparatus than the analysis of standard data. A suitable formal framework is by no means obvious. [^8] Many approaches proposed for analyzing poset-valued data rely on distance measures, see, e.g., Critchlow (1985); Brandenburg et al. (2012). For further discussion see Blocher et al. (2022), Section 2. Fortunately, it turns out that the concept of (data) depth function and the data representation given by formal concept analysis can be promisingly combined and applied to poset-valued random variables. [^9] The representation of posets via formal concept analysis can be expressed as a closure operator, see Definition 1. Therefore, formal concept analysis is not necessary to follow this paper. For more information and discussion of this representation, see Blocher et al. (2022). In this paper, we adapt the concept of depth to poset-valued random variables by introducing the union-free generic (ufg) depth. In general, depth functions define a notion of centrality and outlyingness of observations with respect to an entire data cloud or an underlying distribution, see e.g. Zuo and Serfling (2000); Mosler (2002). So far, depth functions have mainly been applied to $\mathbb{R}^{d}$-valued random variables. Exceptions are e.g. the definition of depth functions on lattices in Schollmeyer (2017b, a), on total orders, see Goibert et al. (2022), and a general discussion on depth functions based on formal concept analysis in Blocher and Schollmeyer (2023). Since we transfer this depth concept to poset-valued data, some classical descriptive statistics can naturally be adapted to this particular non-standard data type.
@@ 1.2 Comparing Machine Learning Algorithms
(I) First, we compare algorithms not with respect to one unidimensional criterion like, e.g., balanced accuracy, but instead look at a whole set of performance measures. We then judge one algorithm as at least as good as another one if it is not outperformed with respect to any of these performance measures. With this, for every data set, we get a partial order of algorithms and since we are not looking at only one, but a whole population/sample of data sets, we get a poset-valued random variable. Importantly, we do not hold the view that there is an underlying true (random) total order together with a coarsening mechanism that generates the (random) partial order. Such views are often termed epistemic, cf., Couso and Dubois (2014). Applications of this view in the context of partial order data can be found for example in Lebanon and Mao (2008); Nakamura et al. (2019). In contrast, within the nomenclature of Couso and Dubois (2014), we see our approach more in the spirit of the ontic view that is usually applied to set-valued data. [^10] Concrete applications can be found e.g. in Plass et al. (2015b) for the ontic, and in Plass et al. (2015a) for the epistemic view. A case where the ontic and the epistemic view coincide is discussed in Schollmeyer (2019). Beyond this, in the field of partial identification in the context of generalizing confidence intervals to confidence regions for the so-called identified set, the question about an ontic vs. an epistemic view is also implicitly asked (without reference to these terms) by asking if such a confidence set should cover the true parameter or the whole identified set with prespecified probability, cf., Stoye (2009). However, since in our case the random objects are partial orders, the term ontic seems to fit not perfectly. Instead, we understand poset-valued data as a special type of non-standard data. [^11] Note that we do not want to generally rule out epistemic treatment, but this is not the focus of this paper. Such a treatment could use that every partial order can be described by the set of all its linear extensions. For further discussions, cf., Blocher et al. (2022).
@@ 1.3 Outline of the Paper
This paper is an extended version of Blocher et al. (2023) presented at the 13th International Symposium on Imprecise Probabilities: Theories and Applications (ISIPTA 2023) conference in Oviedo, Spain. Compared to the ISIPTA 2023 conference, this paper focuses more on the application. We added an entire section on comparison with other methods for evaluating poset-valued data and a sensitivity analysis with respect to the number of performance measures, see Section 7. We also extend the analysis and implementation of the ufg depth, see Sections 5 and 6.
@@
Our paper is organized as follows: We begin with a brief practitioner’s summary, focusing on the special case of classifier comparisons. In Section 3, we briefly discuss the required mathematical definitions and concepts. We give a formal definition of our depth function, the ufg depth, in Section 4 and discuss some of its properties in Section 5. While Section 6 prepares our application by providing implementation details, Section 7 is devoted to applying our framework to specific examples, namely the analysis of the goodness of classification algorithms on different data sets. First, we focus on comparing our benchmarking approach to further benchmarking concepts, see Section 7.1 and afterward, we give a detailed demonstration of the ufg method, see Section 7.2. Section 8 concludes by elaborating on some promising perspectives for future research.
2 Summary for Practitioners
In this section we demonstrate how our general methodology can be used for the problem setting of selecting a reasonable classifier among a set of candidate classifiers within an application situation. Selection methods for classification algorithms are of immense importance for applications in a wide range of industries. Depending on the concrete branch, large amounts of money or even human lives can depend on it, e.g., treatment decision. As many practitioners have broad expertise in their specialty but are less familiar with statistical methodology, reliable guidance on classifier selection is crucial. To qualify as reliable, a selection method should at least meet the following standards:
  • (I)
    Option to make no decision: If the inherent uncertainty of the comparison situation is too large to identify a clear best classifier, the method should report and not obscure this fact.
  • (II)
    Utilization of available information: If the evidence is sufficient to identify a clear best classifier, the method should recognize this and not make overly cautious recommendations.
  • (III)
    Intuitive accessibility: The method should enable the user to develop an intuitive understanding of it (an exact understanding of the underlying mathematics is not necessary).
We present a framework for comparing classifiers that addresses requirements (I) to (III) and is based on two simple observations: First, the suitability of a classifier depends on the specific question, i.e., the specific data set being analyzed. Accordingly, the search should not (solely) be for the best classifier per se, but represent how central/typical a data set and the induced classifier ranking is. Second, the performance of certain classifiers might be incomparable already on one data set because, for example, multiple or vaguely defined performance measures are considered that conflict each other in the quality dimensions.
These two simple observations serve as a guiding principle of our construction: To compare data sets in terms of their centrality (think of the empirical median as most central one and the high and low quantiles as more outlying ones), we use the statistical theory of data depth: the deeper a data set lies (in some abstract space), the more central it is. To account for the complexity of the (potentially multidimensional or vaguely defined) concept of performance, we allow classifiers to be only partially ordered (already when considering only a single data set).
Our resulting framework does indeed manage to meet the above standards: By using partial orders rather than total ones, in the spirit of (I), a clear best classifier is identified only if the underlying evidence is strong enough and always accompanied by information on the typicality of its superiority. In the spirit of (II), our framework avoids the incomparability of classifiers if the evidence is strong enough for a clear favorite (who is still accompanied by information on the typicality of its superiority). Finally, standard (III) is met as well: Since data depth is a generalization of the elementary statistical concept of a quantile (where, e.g., the deepest point corresponds to the median), every user is directly guided by natural intuition when applying the proposed concepts to real-world applications.
In a nutshell, our contributions can be summarized as follows: We offer practitioners the opportunity to reliably compare classification algorithms with regard to several performance measures and taking into account the typicality/centrality of the specific situation analyzed. Beyond the theoretical soundness of our framework, we provide well-documented implementations of our method, which are easily transferable to comparable applications. Last but not least, we propose a customized depth function for our application, the ufg depth, and enable the reliable comparison of classifiers by performing the following simple three-step scheme:
  1. 1.
    Compute the ufg depth function. We provide an R-script that can be easily adapted. The script can be found on GitHub, see https://github.com/hannahblo/Comparing_Algorithms_Using_UFG_Depth.
  2. 2.
    Use the partial order among classifiers corresponding to the highest depth value as a reference and the undominated algorithms as recommendations. We perform this analysis in Section 7.2 on Figure 4.
  3. 3.
    Check how stable the dominance structure is by checking what the partial orders with the $k$ highest depth values have in common. An example analysis using the interesection of the deepest posets can be found in the Section 7.2 discussing Figure 5.
Finally, we want to emphasize that our concept can be used to compare any objects that arise from a sample of instances and where the comparison is based on multiple numeric measures. Thus, the concept introduced here is not limited to classification problems.
@@ 3 Preliminaries
The concept for analyzing poset-valued observations presented in this paper is based on a closure operator on $\mathcal{P}$, see Section 4. In general, a closure operator $\gamma_{\Omega}:2^{\Omega}\to 2^{\Omega}$ on set ${\Omega}$ is an operator which is extensive (for $A\subseteq{\Omega},\>A\subseteq\gamma_{\Omega}(A)$ holds), isotone, (for $A,B\subseteq{\Omega}$ with $A\subseteq B,\>\gamma_{\Omega}(A)\subseteq\gamma_{\Omega}(B)$ is true) and idempotent (for $A\subseteq{\Omega},\;\gamma_{\Omega}(A)=\gamma_{\Omega}(\gamma_{\Omega}(A))$ holds). The set $\gamma_{\Omega}(2^{\Omega})$ is called the closure system and we say that $\gamma(B)$ for $B\subseteq\Omega$ is the closure of $B$. In what follows, we use that every closure operator (and therefore the closure system) can be uniquely described by an implicational system. Generally, an implicational system $\mathcal{I}$ on ${\Omega}$ is defined by a subset of $2^{\Omega}\times 2^{\Omega}$. The implicational system corresponding to the closure operator $\gamma_{\Omega}$ is defined by all pairs $(A,B)\in 2^{\Omega}\times 2^{\Omega}$ satisfying $\gamma_{\Omega}(A)\supseteq\gamma_{\Omega}(B)$. For short, we denote this by $A\to B$. For more details on closure operators and implicational systems, see Bertet et al. (2018).
We aim to define a centrality and outlyingness measure on the set of all posets $\mathcal{P}$ based on a fixed and finite set $M$. In general, functions that measure the centrality of a point with respect to an entire data cloud or an underlying distribution are called (data) depth functions. Depth functions on $\mathbb{R}^{d}$ have been studied intensively by, e.g., Zuo and Serfling (2000) and Mosler and Mozharovskyi (2022), and various notions of depth have been defined, such as Tukey’s depth, see Tukey (1975), and simplicial depth, see Liu (1990). The idea behind the ufg depth introduced here is an adaptation of the simplicial depth on $\mathbb{R}^{d}$ to posets. The simplicial depth on $\mathbb{R}^{d}$ is based on the convex hull/closure operator that is defined as follows:
@@
where $X_{1},\ldots,X_{d+1}\overset{i.i.d.}{\sim}\nu$ with i.i.d. being independent and identically distributed for short. When we consider a sample $x_{1},\ldots,x_{n}\in\mathbb{R}^{d};\>n\in\mathbb{N}$, we use the empirical probability measure instead of a probability measure $\nu$. Thus, for a sample $x_{1},\ldots,x_{n}\in\mathbb{R}^{d}$ with empirical measure $\nu_{n}$ we obtain as empirical simplicial depth
$D_{n}\colon\begin{array}[]{l}\mathbb{R}^{d}\to[0,1],\\ x\mapsto\binom{n}{d+1}\sum_{1\leq i_{1}<\ldots<i_{d+1}\leq n}\mathds{1}_{\gamma_{\mathbb{R}^{d}}\{x_{i_{1}},\ldots,x_{i_{d+1}}\}}(x).\end{array}$
@@ 4 Union-Free Generic Depth on Posets
Now, we introduce the union-free generic (ufg) depth function for posets which is in the spirit of the simplicial depth function, see Section 3. To define the depth function, we start, similar to the simplicial depth, by defining a closure operator on $\mathcal{P}$. This closure operator maps a set of posets $P$ onto the set of posets where each poset is a superset of the intersection of $P$ and a subset of the union of $P$. In other words, any $p$ lying in the closure of $\gamma(P)$ satisfies the following condition: First, every pair $(y_{1},y_{2})\in M\times M$ that lies in every poset in $P$ is also contained in $p$, and second, for every pair $(y_{1},y_{2})$ that lies in $p$, there exists at least one $\tilde{p}\in P$ such that $(y_{1},y_{2})\in\tilde{p}$. Note that while the intersection of posets defines a poset again, this does not hold for the union.
@@
Analogously to the definition of the simplicial depth, we now only consider a subset of $2^{\mathcal{P}}$. We denote this family by $\mathscr{S}$. $\mathscr{S}$ is a proper subset of $2^{\mathcal{P}}$, see Theorem 3 for details, which reduces $2^{\mathcal{P}}$ by redundant elements in the following sense: First, all subsets $P\subseteq\mathcal{P}$ with $\gamma(P)=P$ are trivial and therefore not included in $\mathscr{S}$. Second, if there exists a proper subset $\tilde{P}\subsetneq P$ with $\gamma(\tilde{P})=\gamma(P)$, then $P$ is also not in $\mathscr{S}$. These two properties can be generalized to arbitrary closure systems, and referring to Bastide et al. (2000), we call a set fulfilling these two properties generic. The final reduction is to delete also all sets $P$ where $P$ can be decomposed by a family of proper subsets $(A_{i})_{i\in\{1,\ldots,\ell\}}$ of $P$. Moreover, the union of $(\gamma(A_{i}))_{i\in\{1,\ldots,\ell\}}$ needs to equal $\gamma(P)$. Note that due to isotonicity, the assumption $\cup_{i\in\{1,\ldots,\ell\}}\gamma(A_{i})\subseteq\gamma(P)$ is always true. We call sets respecting this third part union-free. Thus $\mathscr{S}$ consists of elements that are union-free and generic. The following definition summarizes these conditions.
@@
The empirical version of the ufg depth uses the empirical probability measure $\nu_{n}$ given by a sample of posets $\underline{p}=(p_{1},\ldots,p_{n}),\>n\in\mathbb{N}$ instead of the probability measure $\nu$ in Definition 3. Thus, the empirical ufg depth of a poset $p$ is therefore the normalized weighted sum of drawn sets $S\in\mathscr{S}$ which imply $p$.
@@
Note that when restricting $\mathscr{S}$ to the set $\mathscr{S}_{obs}:=\{S\in\mathscr{S}\mid\mbox{ all }p\in S\mbox{ are observed}\}$, this does not change the depth value. This holds since for other elements $S\in\mathscr{S}$, the empirical measure for at least one $p\in S$ is zero.
@@
Returning to Example 1, suppose that we observe $(p_{1},p_{2},p_{3})$. Then for the trivial poset $p_{\Delta}$, the empirical ufg depth is $D_{n}(p_{\Delta})=1/2$. For the set $p_{4}$ given by the transitive hull of $\{(y_{3},y_{1})\}$, the value of the empirical ufg depth is zero. For $p_{total}$ given by the transitive hull of $\{(y_{1},y_{3}),(y_{3},y_{2})\}$, the empirical ufg depth value is again zero.